OpenStudy (anonymous):

Meghan has 100 dollars more than Andrew has. After Meghan spends 20 dollars on groceries, she will have five times as much money as Andrew. How much money does Andrew have?

OpenStudy (anonymous):

do i have to duble the cost since its say five times or multiply 5

OpenStudy (anonymous):

multiply by 5

OpenStudy (anonymous):

i got 120 is that correct

OpenStudy (anonymous):

Meghan = x+100 Andre = x (x+100 - 20) = 5x x + 80 = 5x 80 = 5x - x 80 = 4x x = 80/4 x = 20 so Andre have 20 dollar

OpenStudy (anonymous):

i think the answer is 20 dollar for Andrew

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

ur welcome .., am i right??